One number exceeds another by ten and their sum is 34. Find the larger number.
Answer by Earlsdon(6294)

Let x = the larger number and y = the smaller number.
x+y = 34 and...
x = y+10 Substitute this into the eqution above.
(y+10)+y = 34 Simplify.
2y+10 = 34 Subtract 10 from both sides.
2y = 24 Divide both sides by 2.
y = 12 and...
x = 12+10
x = 22.
The larger number is 22.